Caro Manufacturing has two production departments, Machining and Assembly, and two service departments, Maintenance and Cafeteria. Direct costs for each department and the proportion of service costs used by the various departments for the month of August follow:

Proportion of Services Used by

Department

Direct Costs

Maintenance

Cafeteria

Machining

Assembly

Machining

$

96,000

Assembly

70,400

Maintenance

43,600

0.2

0.5

0.3

Cafeteria

35,000

0.6

0.2

0.2

Compute the allocation of service department costs to producing departments using the direct method. (Do not round intermediate calculations.)

Matching =

Assembly=
Answer #1

Answer is

Machining = $44750

Assembly:=$33850

Cost alloctiated in the ratio of 5:3 and 2:2

Department Direct Costs Machining Assembly
Maintenance 5:3 43,600 27250 16350
Cafeteria 2:2 35,000 17500 17500
Cost Allocated$ $44750 $33850
Department Direct Costs Machining Assembly
Maintenance 5:3 43600 =43600*5/8=27250 =3/8*43600=16350
Cafeteria 2:2 35000 =35000*2/4=17500 =2/4*35000=17500